OpenStudy (anonymous):

2/7x = 10 (explain if possible)

OpenStudy (anonymous):

You need to isolate the variable (7x) first

OpenStudy (anonymous):

ok how do I do that since its attached to the 2 as well?

OpenStudy (anonymous):

What happens if you multiply each side of the equation by 1/2.

OpenStudy (anonymous):

so I multiple 2/7 by 10/1?

OpenStudy (anonymous):

Sorry connection issues...What is 2/7x times 1/2?

OpenStudy (anonymous):

where are you getting 1/2 from???

OpenStudy (anonymous):

Try moving the 7x to the right hand side of the equation.

OpenStudy (anonymous):

\(2 = (7x)(10)\)

OpenStudy (anonymous):

Can you solve for x now?

OpenStudy (anonymous):

Forget everything I said ..easy way to solve. Multiply both sides by 7/2.

OpenStudy (anonymous):

so if I minus 10 I would then end up with -8 = 7x now if I divide I get a decimial and that's def wrong?

OpenStudy (anonymous):

You can't minus here. @c1282

OpenStudy (anonymous):

You have to do multiplication.

OpenStudy (anonymous):

so then how are we isolating x?the answer cant b -8= 7x

OpenStudy (anonymous):

The 10 and the 7x.

OpenStudy (anonymous):

m not trying to get the answer but I do not understand with 10 and 7x what im suppose to do can u pls try breaking it down even if u use different numbers

OpenStudy (anonymous):

nevermind ill see if I find another way thank u

OpenStudy (anonymous):

Okay. \(\dfrac{2}{7x} = 10\) We want to solve for x- we don't want to leave it in a fraction form. Now 10 is on the other side of the = sign. Here we can use 'cross multiplication' between the 7x and 10. So we can move the 7x to the right hand side, but its multiplied with the 10. \(2 = (10)(7x)\) Are you getting it so far?

OpenStudy (anonymous):

\[\frac{ 2 }{ 7x } x \frac{ 7 }{ 2 } = x\]

OpenStudy (anonymous):

Now do the same for the other side of the equation\[10 x \frac{ 7 }{ 2}\]